OpenStudy (anonymous):

solve the equation. show your work 3x=45

OpenStudy (gorv):

divide both side by 3

OpenStudy (anonymous):

Okay, so for this, you first need to isolate the x and to do this you need to remove the 3 that is multiplying with it. So you divide 3 to by each side to cancel out the multiplication. 3x = 45 x/3 =45/3 x= 15

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

You could also check your answer by adding in 15 in the spot of the variable x. 3x = 45 3x 15 = 45

OpenStudy (anonymous):

No no questions. Thank you that is what I was trying to understand! :)

OpenStudy (anonymous):

divide 3x=45

OpenStudy (anonymous):

Thanks gorv and mcnabkt, I also needed an example to understand the equation

OpenStudy (anonymous):

One thing, I forgot to write something. 3x = 45 x/3 =45/3 ---> That should be written as 3x/3 because that is technically what you're doing. x= 15

OpenStudy (anonymous):

easy. Just divide both sides by 3